A production function has the following form Q = 4KL. Price of labor is PL = 8 and price of capital is PK = 4. How many units of labor and capital the firm hires, if it minimizes its costs and its goal is to produce 800 units of output? Calculate total costs.


1
Expert's answer
2021-11-08T09:10:34-0500

Solution:

Cost minimization is where: MPL/MPK = w/r or MPL/w = MPK/r

Q = 4KL

MPL = "\frac{\partial Q} {\partial L} = 4K"


MPK = "\frac{\partial Q} {\partial K} = 4L"

w = 8

r = 4

"\frac{4K} {4L} = \frac{8} {4}"

K = 2L

Substitute in the production function:

Q = 4KL

800 = 4(2L) L

800 = 8L2

L2 = 100

L = 10

K = 2L = 2"\times"10 = 20

K = 20

The firm will hire 10 units of labor and 20 units of capital.

 

TC = wL + rK

TC = 8(10) + 4(20) = 80 + 80 = 160

Total Costs = 160